A 240 cm length of string has a mass of 4.3 g. It is stretched with a tension of 8.0 N between fixed supports. (a) What is the wave speed for this string? (b) What is the lowest resonant frequency of this string?
Answer #1

a)  linear mass = [0.0043 kg / 2.40 m]

V = sqrt ( 8.0 N / [0.0043 kg / 2.40 m] )

wave speed for this string = 66.82 m/s

--------------------------------------------------

b) L = (1/2) (V/f)

f = V / 2L = 66.82 / (2 * 2.40 m)

lowest resonant frequency of this string = 13.92 Hz